Grade 12Physical Chemistry

How to calculate equivalent weight in a diproportionation redox reaction? Give an example.

Anurag Kshatri
13 Years agoGrade 12

1 Answer

Aman  Bansal
13 Years ago

Dear Anurag,

Equivalent weight of Oxidising and Reducing Agents


Equivalent weight of a substance (oxidant or reductant) is equal to molecular weight divided by number of electrons lost or gained by one molecule of the substance in a redox reaction.

Eq. wt. of O. A. 

                       =

Eq. wt. of R. A.  = 

                       = 

  Table : Equivalent weight of few oxidising/reducing agents
 

Agents

O. N.

Product

O. N.

Change in
O. N. per
atom

Total 
Change in
O. N. per  mole

Eq. wt.

+ 6

+ 3

3

3 × 2 = 6

Mol. wt./6

+ 3

+ 4

1

1 × 2 = 2

Mol. wt./2

+ 2

+ 2.5

0.5

0.5 × 2 = 1

Mol. wt./1

– 1

– 2

1

1 × 2 = 2

Mol. wt./2

– 1

0

1

1 × 2 = 2

Mol. wt./2


(Acidic medium)

+ 7

+ 2

5

5 × 1 = 5

Mol. wt./5


(Neutral medium)

+ 7

+ 4

3

3 × 1 = 3

Mol. wt./3


(Alkaline medium)

+ 7

+ 6

1

1 × 1 = 1

Mol. wt./1


Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

https://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert